# Latent Multi-Head Attention for Small Language Models

### v1

@@ Abstract.
We present the first comprehensive study of latent multi-head attention (MLA) for small language models, revealing interesting efficiency-quality trade-offs. Training  30M-parameter GPT models on 100,000 synthetic stories, we benchmark three architectural variants: standard multi-head attention (MHA), MLA, and MLA with rotary positional embeddings (MLA+RoPE). Our key finding is that MLA+RoPE with half-rank latent dimensions ($r=d/2$) achieves a 45% KV-cache memory reduction while incurring only a 0.3% increase in validation loss (essentially matching MHA’s quality)—a Pareto improvement for memory-constrained deployment. We further show that RoPE is crucial for MLA in small models: without it, MLA underperforms vanilla attention by 3-5%, but with RoPE, it surpasses vanilla by 2%. Inference benchmarks on NVIDIA A100 GPUs reveal that MLA with $r=d/2$ achieves a 1.4× speedup over full-rank MLA while maintaining the memory savings. GPT-4 evaluations corroborate perplexity results, with MLA+RoPE achieving the highest quality scores (7.4/10) across grammar, creativity, and consistency metrics. Code and models will be released upon acceptance.
@@ 1. Introduction
The deployment of language models in resource-constrained environments—edge devices, embedded systems, and cost-sensitive applications—necessitates rethinking the efficiency-capability trade-off (wang2024slm, ). While large models dominate benchmarks (brown2020language, ; openai2023gpt4, ), recent work on TinyStories (eldan2023tinystories, ) demonstrates that models under 10M parameters can achieve linguistic fluency in constrained domains. This suggests that architectural efficiency, not just parameter count, may be key to practical deployment.
We investigate two complementary innovations: multi-head latent attention (MLA) (meng2025transmla, ), which compresses key-value (KV) representations through learned low-rank projections, and rotary positional embeddings (RoPE) (su2021roformer, ), which encode relative positions without explicit embedding parameters. While MLA has shown promise in large models like DeepSeek-V2 (deepseek2024v2, ), its behavior in small-scale regimes remains unexplored.
@@
  • Q1: Can MLA’s memory efficiency benefits transfer to 30M-parameter models without significant quality loss?
  • Q2: How does the latent dimension $r$ affect the memory-latency- quality trade-off surface?
  • Q3: Do automated (GPT-4) evaluations corroborate perplexity differences between architectures?
@@
    GPT-4 evaluations align with perplexity: Automated quality scores strongly correlate with validation loss, with MLA+RoPE achieving the highest ratings.
2. Background and Method
@@ 2.1. Multi-Head Latent Attention
where $X\in\mathbb{R}^{n\times d}$ represents input embeddings and $W_{h}^{Q,K,V}\in\mathbb{R}^{d\times d_{k}}$ are projection matrices with $d_{k}=d/H$ being the per-head dimension.
MLA introduces a bottleneck by factorizing the key and value projections:
@@
where $W_{h}^{K\downarrow}\in\mathbb{R}^{d\times r}$, $W_{h}^{K\uparrow}\in\mathbb{R}^{r\times d_{k}}$, and $r<d_{k}$ is the latent dimension. During inference, only $K_{h}^{\text{latent}},V_{h}^{\text{latent}}\in\mathbb{R}^{n\times r}$ are cached, reducing memory from $O(nHd_{k})$ to $O(nHr)$. Thus, $r=d_{k}$ yields no compression (MLA becomes equivalent to standard MHA), whereas smaller $r$ compresses the KV cache proportionally. In our experiments, we evaluate $r\in\{d_{k},\,d_{k}/2,\,d_{k}/4\}$ to quantify these trade-offs.
@@ 2.3. Training Setup
Table 1. Model configurations (GPT-2 tokenizer, tied soft-max).
@@ 3.1. Quality-Efficiency Trade-off
Table 2 presents our core findings on the 9L-512d configuration, including extreme compression ratios up to $r=d/32$. MLA with full rank ($r=d$, no compression) underperforms vanilla MHA by 3.2% without RoPE, confirming that even without compression, the additional projections can harm small models. However, adding RoPE not only recovers this loss but achieves a 2.1% improvement. Most surprisingly, MLA+RoPE with half-rank compression ($r=d/2$) nearly matches vanilla MHA quality (only 0.3% degradation) while providing substantial efficiency gains.
Our investigation of extreme compression ratios reveals a critical phase transition around $r=d/16$. The validation loss increases follow an approximately exponential pattern beyond $r=d/8$, with the model becoming effectively unusable at $r=d/32$ where loss increases by over 20% even with RoPE. This aligns with theoretical predictions from information bottleneck theory and empirical observations in larger models, though small models show heightened sensitivity to extreme compression.
@@
The critical observation is that beyond $r=d/8$, both MLA variants exhibit catastrophic degradation. Models with $r=d/16$ produce semi-coherent but highly repetitive text, while $r=d/32$ configurations generate fairly random token sequences. This phase transition occurs more abruptly in small models compared to the gradual degradation reported in large-scale implementations like DeepSeek-V2, suggesting fundamental differences in how model capacity affects compressibility.
@@ 3.2. GPT-4 Quality Evaluation
The GPT-4 evaluations strongly corroborate our perplexity findings. MLA+RoPE ($r=d/2$) achieves the highest scores across all dimensions, with particularly notable improvements in consistency (+1.7 over MHA) and creativity (+1.3). The pure MLA model shows significant degradation, especially in consistency (-0.8 compared to MHA), aligning with our observation that positional information is crucial for maintaining coherent narratives with compressed attention.
@@ 3.3. Qualitative Analysis
We performed detailed qualitative comparisons of stories generated by the 12L-1024d models, revealing distinct failure modes for each architecture. Each model was prompted with identical story beginnings and asked to generate token continuations.
@@
Vanilla MHA outputs demonstrated better grammatical fluency and maintained more consistent character identities throughout the stories. However, these stories often suffered from predictable plot developments and occasionally ended abruptly without proper narrative resolution. While technically correct, they tended toward safe, formulaic story structures reminiscent of the most common patterns in the training data.
In contrast, MLA+RoPE consistently produced the most coherent and engaging narratives. The model demonstrated superior ability to track narrative state across longer contexts, rarely losing track of established story elements or character relationships.
These qualitative observations align closely with the quantitative GPT-4 evaluations. The MLA+RoPE model’s superior narrative consistency (+30% over vanilla MHA) reflects its enhanced ability to maintain story coherence through the combination of compressed attention and relative position encoding. Meanwhile, MLA without RoPE’s low consistency score (4.8/10) directly corresponds to the frequent coherence failures observed in manual analysis.
The differences were most pronounced in longer stories, where maintaining narrative coherence becomes increasingly challenging. For very short continuations, all models produced acceptable outputs, suggesting that the architectural differences primarily manifest when sustained context tracking is required.
@@ 3.4. Inference Performance Analysis
Figure 1 presents a detailed breakdown of token–level throughput and KV‑cache memory footprint on a single NVIDIA A100 80GB. We also use larger batch sizes of $32$ sequences, which is a common setting for server–side batched decoding workloads.
@@
The left panel of Figure 1 (a) confirms that vanilla MHA saturates early: tokens/sec decrease from $210$ tok/s at batch 1 to $170$ tok/s at batch 32 because attention bandwidth becomes memory–bound once the hardware queues are full. In contrast, all MLA variants improve as the batch grows thanks to the smaller $r\times r$ attention matrix that fits in on‑chip SRAM, leading to better L2 cache locality. With $r=d/4$, throughput climbs from $145\!\to\!186$ tok/s and only drops below MHA up to batch 16; at batch 32 it overtakes MHA by 9%. This confirms that MLA can close the gap to dense attention with sufficient parallelism.
@@
Adding RoPE increases arithmetic intensity, so MLA+RoPE trails its no‑RoPE counterpart by $\approx\!15$ tok/s across the board. Yet MLA+RoPE with $r=d/2$ still reaches $97$ tok/s at batch 32—1.4$\times$ faster than MLA+RoPE with $r=d$—highlighting that positional encodings and latent compression are largely orthogonal knobs.
@@
The right panel (b) reproduces the KV‑cache memory trend from Figure 1 (b): memory grows linearly with latent dimension, matching the theoretical $O(r)$ prediction. Even at the most aggressive compression ($r=d/8$) we observe a 7.5$\times$ reduction relative to MHA, pushing per‑token cache usage below $4$ kB.
@@
MLA with $r=d/4$ is an optimal spot for small models: it delivers (1) near‑MHA throughput for interactive workloads (~0.9$\times$ at batch 1), (2) higher throughput than MHA for batched serving, and (3) up to 6$\times$ KV‑cache savings, enabling longer context windows on memory‑constrained GPUs.
@@
Figure 1. Inference performance trade‑offs. (a) MLA with $r=d/4$ overtakes MHA once batching is large enough, achieving a 9% throughput win at batch 32. (b) Memory savings scale linearly with latent dimension, closely matching theoretical $O(r)$ behavior.
4. Related Work
Efficient Attention Mechanisms.
Beyond MLA, several approaches address attention’s computational bottlenecks. Grouped-Query Attention (GQA) (ainslie2023gqa, ) shares KV projections across head groups, while Multi-Query Attention (MQA) (shazeer2019fast, ) uses a single shared KV. FlashAttention (dao2022flashattention, ) optimizes memory access patterns without changing attention’s mathematical formulation. Our work shows MLA provides orthogonal benefits that compound with these optimizations.
Small Language Models.
Recent work challenges the ”bigger is better” paradigm. MobiLLM (mobilellm2024, ) demonstrates sub-billion parameter models can match larger ones through careful design. MiniCPM (minicpm2024, ) achieves GPT-3.5 performance at 2.3B parameters via architectural search. Our findings align with this trend, showing that innovation at small scale yields significant returns.
Position Encodings.
While RoPE has become standard in large models (touvron2023llama2, ), its interaction with architectural choices remains understudied. ALiBi (press2022alibi, ) provides an alternative through attention biasing, while recent work explores learned continuous encodings (golovneva2024contextual, ). Our analysis reveals position encoding choice critically affects compressed attention methods.
@@ 5. Limitations and Future Work
While our results demonstrate promising efficiency gains, several limitations warrant careful consideration:
  • Domain specificity: Our experiments exclusively use the TinyStories dataset, which contains simple children’s narratives. The effectiveness of MLA+RoPE on more complex domains (technical documentation, code, multilingual text) remains unexplored. The constrained vocabulary and narrative structure may not reveal failure modes that emerge in diverse text generation tasks.
@@
  • Fixed compression ratios: We evaluated only three compression levels ($r\in\{d,d/2,d/4\}$). Adaptive compression schemes that vary $r$ across layers or heads could potentially achieve better efficiency-quality trade-offs. Additionally, learned compression ratios might automatically discover optimal configurations.
  • Hardware-specific results: Our inference benchmarks target NVIDIA A100 GPUs. Performance characteristics may differ substantially on other accelerators (TPUs, Apple Silicon) or CPUs. Memory access patterns and computational intensity vary across platforms, potentially changing the relative advantages of each approach.
  • Evaluation methodology: While GPT-4 evaluations provide scalable quality assessment, they may not capture all aspects of generation quality important to end users. Human evaluation on diverse tasks would provide more comprehensive validation.
@@ 6. Conclusion
This work demonstrates that multi-head latent attention, when combined with rotary position embeddings, provides a Pareto improvement for small language models: 45% memory reduction and 1.4× inference speedup with minimal quality loss. Our key insight—that RoPE is essential for maintaining quality with compressed attention in small models—has immediate practical implications for deployment in memory-constrained environments.
More broadly, our results challenge the prevailing narrative that model scale dominates all other considerations. For the millions of potential deployments where every megabyte and millisecond matter—from mobile devices to embedded systems—architectural innovation remains paramount. As foundation models democratize, the ability to run capable models on commodity hardware will determine their real-world impact. MLA+RoPE represents one step toward this goal, demonstrating that careful co-design of attention mechanisms and position encodings can unlock significant efficiency gains without sacrificing quality.
We hope this work inspires further research into efficient architectures that bring capable AI to the edge, making language models accessible beyond data centers and high-end hardware.
@@ Acknowledgments
Appendix A Full Experimental Results
Table 4 presents complete results across all 8 model configurations and 3 architectures.
@@ Appendix A Full Experimental Results
Appendix B Extended GPT-4 Evaluation Results
Table 5 presents comprehensive GPT-4 evaluation scores across all compression ratios for the 12L-1024d configuration. The scores reveal distinct degradation patterns for each quality metric as compression increases.

### v2

@@ Abstract.
We present the first comprehensive study of latent multi-head attention (MLA) for small language models, revealing interesting efficiency-quality trade-offs. Training  30M-parameter Generative Pre-trained Transformer (GPT) models on 100,000 synthetic stories, we benchmark three architectural variants: standard multi-head attention (MHA), MLA, and MLA with rotary positional embeddings (MLA+RoPE). Our key finding is that MLA+RoPE with half-rank latent dimensions ($r=d/2$) achieves a 45% Key-Value (KV)-cache memory reduction while incurring only a 0.3% increase in validation loss (essentially matching MHA’s quality), a Pareto improvement for memory-constrained deployment. We further show that RoPE is crucial for MLA in small models: Without it, MLA underperforms vanilla attention by 3-5%, but with RoPE, it surpasses vanilla by 2%. Inference benchmarks on NVIDIA A100 Graphics Processing Units (GPUs) reveal that MLA with $r=d/2$ achieves a 1.4× speedup over full-rank MLA while maintaining memory savings. GPT-4 evaluations corroborate perplexity results, with MLA+RoPE achieving the highest quality scores (7.4/10) on grammar, creativity, and consistency metrics. The code and models will be released upon acceptance.
@@ 1. Introduction
The deployment of language models in resource-constrained environments, such as edge devices, embedded systems, and cost-sensitive applications, requires rethinking the efficiency-capability trade-off (wang2024slm, ). Although large models dominate benchmarks (brown2020language, ; openai2023gpt4, ), recent work on TinyStories (eldan2023tinystories, ) demonstrates that models with 10M parameters can achieve linguistic fluency in constrained domains. This suggests that architectural efficiency, not just the parameter count, may be a key to practical deployment.
We investigate two complementary innovations: multi-head latent attention (MLA) (meng2025transmla, ), which compresses key-value (KV) representations through learned low-rank projections, and rotary positional embeddings (RoPE) (su2021roformer, ), which encode relative positions without explicit embedding parameters. Although MLA has shown promise in large models such as DeepSeek-Version2(V2) (deepseek2024v2, ), its behavior in small-scale regimes remains unexplored.
@@
  • Question 1 (Q1): Can MLA’s memory efficiency benefits transfer to 30M-parameter models without significant quality loss?
  • Question 2 (Q2): How does the latent dimension $r$ affect the quality of the memory-latency trade-off surface?
  • Question 3 (Q3) Do automated evaluations (GPT-4) corroborate differences in perplexity between architectures?
@@
    GPT-4 evaluations align with perplexity: Automated quality scores are strongly correlated with validation loss, with MLA+RoPE achieving the highest scores.
2. Background
@@ 2.1. Multi-Head Latent Attention
where $X\in\mathbb{R}^{n\times d}$ represents the input embeddings and $W_{h}^{Q,K,V}\in\mathbb{R}^{d\times d_{k}}$ are projection matrices with $d_{k}=d/H$ being the dimension per head.
MLA introduces a bottleneck by factorizing the key and value projections.
@@
where $W_{h}^{K\downarrow}\in\mathbb{R}^{d\times r}$, $W_{h}^{K\uparrow}\in\mathbb{R}^{r\times d_{k}}$, and $r<d_{k}$ is the latent dimension. During inference, only $K_{h}^{\text{latent}},V_{h}^{\text{latent}}\in\mathbb{R}^{n\times r}$ are cached, reducing memory from $O(nHd_{k})$ to $O(nHr)$. Thus, $r=d_{k}$ yields no compression, as MLA becomes equivalent to standard MHA; smaller $r$ values compress the KV cache proportionally. In our experiments, we evaluate $r\in\{d_{k},\,d_{k}/2,\,d_{k}/4\}$ to quantify these trade-offs.
@@ 2.2. Rotary Position Embeddings
3. Related Work
Efficient Attention Mechanisms.
Beyond MLA, several approaches address attention’s computational bottlenecks. Grouped query attention (GQA) (ainslie2023gqa, ) shares KV projections between head groups, while Multi-Query Attention (MQA) (shazeer2019fast, ) uses a single shared KV. FlashAttention (dao2022flashattention, ) optimizes memory access patterns without changing the mathematical formulation of attention. Our work shows that MLA provides orthogonal benefits that compound with these optimizations.
Small Language Models.
Recent work challenges the ”bigger is better” paradigm. MobiLLM (mobilellm2024, ) demonstrates that subbillion parameter models can match larger ones through careful design. MiniCPM (minicpm2024, ) achieves the performance of GPT-3.5 with the parameters 2.3B through architectural search. Our findings align with this trend, showing that innovation on a small scale yields significant returns.
Position Encodings.
Although RoPE has become standard in large models (touvron2023llama2, ), its interaction with architectural choices remains understudied. ALiBi (press2022alibi, ) provides an alternative through attention biasing, while recent work explores learned continuous encodings (golovneva2024contextual, ). Our analysis reveals that position encoding choice critically affects compressed attention methods.
4. Method
@@ 4.1. Training Setup
Table 1. Model configurations using GPT-2 tokenizer and tied softmax output layer. The models were trained on a corpus of 10,000 books.
@@ 5.1. Quality-Efficiency Trade-off
Table 2 presents our main findings on the 9L-512d configuration, including extreme compression ratios of up to $r=d/32$. MLA with full rank ($r=d$, without compression) underperforms vanilla MHA by 3.2% without RoPE, confirming that even without compression, additional projections can harm small models. However, adding RoPE not only recovers this loss, but achieves a 2.1% improvement. Most surprisingly, MLA+RoPE with half-rank compression ($r=d/2$) almost matches the quality of vanilla MHA (only 0. 3% degradation) while providing substantial efficiency gains.
Our investigation of extreme compression ratios reveals a critical phase transition around $r=d/16$. The validation loss increases follow an approximately exponential pattern beyond $r=d/8$, with the model becoming effectively unusable at $r=d/32$ where the loss increases by over 20% even with RoPE. This aligns with theoretical predictions from information bottleneck theory and empirical observations in larger models, though small models show heightened sensitivity to extreme compression.
@@
The critical observation is that beyond $r=d/8$, both MLA variants exhibit catastrophic degradation. Models with $r=d/16$ produce semicoherent but highly repetitive text, while $r=d/32$ configurations generate fairly random token sequences. This phase transition occurs more abruptly in small models compared to the gradual degradation reported in large-scale implementations like DeepSeek-V2, suggesting fundamental differences in how model capacity affects compressibility.
@@ 5.2. GPT-4 Quality Evaluation
The GPT-4 evaluations strongly corroborate our perplexity findings. MLA+RoPE ($r=d/2$) achieves the highest scores in all dimensions, with particularly notable improvements in consistency (+1.7 over MHA) and creativity (+1.3). The pure MLA model shows significant degradation, especially in consistency (-0.8 compared to MHA), which is consistent with our observation that positional information is crucial to maintaining coherent narratives with compressed attention.
@@ 5.3. Qualitative Analysis
We performed detailed qualitative comparisons of the stories generated by the 12L-1024d models, revealing distinct failure modes for each architecture. Each model was prompted with identical story beginnings and asked to generate token continuations.
@@
Vanilla MHA outputs demonstrated better grammatical fluency and maintained more consistent character identities throughout the stories. However, these stories often suffered from predictable plot developments and occasionally ended abruptly without proper narrative resolution. Although technically correct, they tended toward safe, formulaic story structures reminiscent of the most common patterns in the training data.
In contrast, MLA+RoPE consistently produced the most coherent and engaging narratives. The model demonstrated a superior ability to track the state of the narrative in longer contexts, rarely losing track of established story elements or character relationships.
These qualitative observations align closely with the quantitative evaluations of GPT-4. The superior narrative consistency of the MLA+RoPE model (+30% over vanilla MHA) reflects its enhanced ability to maintain story coherence through the combination of compressed attention and relative position encoding. Meanwhile, MLA without RoPE’s low consistency score (4.8/10) directly corresponds to the frequent coherence failures observed in manual analysis.
The differences were most pronounced in longer stories, where maintaining narrative coherence becomes increasingly challenging. For very short continuations, all models produced acceptable output, suggesting that the architectural differences manifest primarily when sustained context tracking is required.
@@ 5.4. Inference Performance Analysis
Figure 1 presents a detailed breakdown of token level throughput and KV‑cache memory footprint on a single NVIDIA A100 80GB. We also use larger batch sizes of $32$ sequences, which is a common setting for server–side batched decoding workloads.
@@
The left panel of Figure 1 (a) confirms that vanilla MHA saturates early: tokens/sec decrease from $210$ tok/s at batch 1 to $170$ tok/s at batch 32 because attention bandwidth becomes memory–bound once the hardware queues are full. In contrast, all MLA variants improve as the batch grows thanks to the smaller $r\times r$ attention matrix that fits in on‑chip SRAM, leading to better L2 cache locality. With $r=d/4$, throughput increases from $145\!\to\!186$  tac / s and only drops below MHA up to batch 16; at batch 32 it overtakes MHA by 9%. This confirms that MLA can close the gap to dense attention with sufficient parallelism.
@@
Adding RoPE increases the arithmetic intensity, so MLA + RoPE trails its counterpart without RoPE by $\approx\!15$ tok / s throughout the board. However, MLA + RoPE with $r=d/2$ still reaches $97$ tok/s at batch 32—1.4$\times$ faster than MLA+RoPE with $r=d$—highlighting that positional encodings and latent compression are largely orthogonal knobs.
@@
The right panel (b) reproduces the KV‑cache memory trend from Figure 1 (b): memory grows linearly with the latent dimension, matching the theoretical prediction $O(r)$. Even in the most aggressive compression ($r=d/8$) we observe a reduction 7.5$\times$ relative to MHA, pushing the use of the cache per token below $4$  kB.
@@
MLA with $r=d/4$ is an optimal spot for small models: it delivers (1) near‑MHA throughput for interactive workloads (~0.9$\times$ at batch 1), (2) higher throughput than MHA for batched serving, and (3) up to 6$\times$ KV cache savings, allowing longer context windows on memory‑constrained GPUs.
@@
Figure 1. Inference performance trade-offs for MLA and MHA. (a) Throughput of MLA with reduced dimension $r=d/4$ exceeds that of MHA at batch size 32, resulting in a 9% throughput increase. (b) Memory savings of MLA scale linearly with latent dimension, closely matching the expected $O(r)$ complexity.
@@ 6. Limitations and Future Work
Although our results demonstrate promising efficiency gains, several limitations warrant careful consideration.
  • Domain specificity: Our experiments exclusively use the TinyStories dataset, which contains simple children’s narratives. The effectiveness of MLA+RoPE in more complex domains (technical documentation, code, multilingual text) remains unexplored. The constrained vocabulary and narrative structure may not reveal the failure modes that emerge in diverse text generation tasks.
@@
  • Fixed compression ratios: We evaluated only three compression levels ($r\in\{d,d/2,d/4\}$). Adaptive compression schemes that vary $r$ between layers or heads could potentially achieve better efficiency-quality trade-offs. Additionally, learned compression ratios might automatically discover optimal configurations.
  • Hardware-specific results: Our inference benchmarks target NVIDIA A100 GPUs. Performance characteristics may differ substantially on other accelerators (Tensor Processing Units (TPUs), Apple Silicon) or Central Processing Units (CPUs). Memory access patterns and computational intensity vary between platforms, potentially changing the relative advantages of each approach.
  • Evaluation methodology: Although GPT-4 evaluations provide scalable quality assessment, they may not capture all aspects of generation quality important to end users. Human evaluation on various tasks would provide more comprehensive validation.
@@ 7. Conclusion
This work demonstrates that multihead latent attention, when combined with rotary position embeddings, provides a Pareto improvement for small language models: 45% memory reduction and 1.4× inference speedup with minimal quality loss. Our key insight, that RoPE is essential for maintaining quality with compressed attention in small models, has immediate practical implications for deployment in memory-constrained environments.
More broadly, our results challenge the prevailing narrative that the model scale dominates all other considerations. For the millions of potential deployments where every megabyte and millisecond matter, from mobile devices to embedded systems, architectural innovation remains paramount. As foundation models democratize, the ability to run capable models on commodity hardware will determine their real-world impact. MLA+RoPE represents one step toward this goal, demonstrating that careful co-design of attention mechanisms and position encodings can unlock significant efficiency gains without sacrificing quality.
We hope that this work inspires further research into efficient architectures that bring capable Artificial Intelligence (AI) to the edge, making language models accessible beyond data centers and high-end hardware.
@@ Acknowledgments
Appendix A Appendix A: Full Experimental Results
Table 4 presents complete results across the eight model configurations and three architectures.
@@ Appendix A Appendix A: Full Experimental Results
Appendix B Appendix B: Extended GPT-4 Evaluation Results
Table 5 presents comprehensive GPT-4 evaluation scores in all compression ratios for the 12L-1024d configuration. The scores reveal distinct degradation patterns for each quality metric as compression increases.